# A Perspective on AI-Guided Molecular Simulations in VR: Exploring Strategies for Imitation Learning in Hyperdimensional Molecular Systems

### v1

@@ 3.1.1 Manipulation tasks
Manipulation tasks involve robots handling, moving, or altering the state of objects in their environment. In recent years, Imitation learning has been instrumental in teaching robots to perform such tasks with precision and adaptability[31]. For instance, VIOLA [53] a novel IL approach that was implemented and deployed into a real-life robot, outperforms state-of-the-art methods by 45.8% in success rate. This is achieved through the use of a pre-trained vision model which is put into a transformer-like architecture. The authors created a policy to detect task-driven relevant regions for action mapping. Another novel hybrid imitation learning (HIL) framework combines behavior cloning (BC) and state cloning (SC) methods to efficiently learn manipulation tasks like pick-and-place and stacking[16]. This approach has been shown to significantly improve training efficiency and policy flexibility, demonstrating a performance improvement and faster training time compared to pure BC methods. Hua et al. [13] emphasize the efficiency of learning from good samples and the potential for combining reinforcement learning mechanisms to improve the speed and accuracy of imitation learning. They specifically addressed the application of imitation learning in robot manipulation by observing expert demonstrations, which can be generalized to other unseen scenarios.
@@ 3.1.2 Locomotion and navigation
In robotics, locomotion and navigation are two fundamental aspects that enable robots to move and operate within their environments effectively. These concepts are crucial for the development of autonomous systems that can perform a wide range of tasks, from simple delivery services to complex exploration missions. Locomotion refers to the various methods that robots use to move from one place to another. This movement can be achieved through different mechanisms, depending on the robot’s design and the environment it is intended to operate in. Navigation involves the process by which a robot determines its position in the environment and plans a path to reach a specific destination. It encompasses several key competencies: Self-Localization; The ability of a robot to establish its own position and orientation within a frame of reference. Path Planning; Once the robot knows its location, path planning involves determining the most efficient or safest route to reach the desired destination. Map-Building and Map Interpretation; For effective navigation, robots often need to construct or utilize maps of their environment. This involves sensing the surroundings to identify obstacles, pathways, and other relevant features, and interpreting this information to make navigation decisions.
In the research paper "Learning to Walk by Steering: Perceptive Quadrupedal Locomotion in Dynamic Environments," Seo et al. [41] introduce PRELUDE, a hierarchical learning framework designed to enhance the navigation and locomotion capabilities of quadrupedal robots in dynamic and cluttered environments. The framework divides the problem into two levels: high-level navigation decision-making and low-level gait generation. The high-level controller is trained using imitation learning from human demonstrations collected with a steerable cart, enabling the robot to acquire complex navigation behaviors. The low-level gait controller is trained through reinforcement learning, allowing the discovery of versatile gait patterns through trial and error. The effectiveness of PRELUDE is demonstrated through simulations and hardware experiments, showing significant improvements over state-of-the-art reinforcement learning methods in terms of success rate and travel distance in various environmental conditions[41]. This work exemplifies the application of imitation learning in robotics, particularly in the development of autonomous systems capable of agile and adaptive movement in real-world scenarios.
@@ 3.1.3 Human robot interaction
Human interaction tasks in robotics involve robots engaging in various forms of social interaction and cooperation with humans to achieve shared goals. These tasks encompass direct physical interaction, such as assisting with lifting objects or providing physical therapy, as well as collaborative interaction, where robots and humans work together to complete tasks like assembling products on a manufacturing line. Remote interaction, where humans control or collaborate with robots from a distance, also falls under the umbrella of human-robot interaction tasks. Additionally, robots should be able to learn new tasks from human demonstrations and proactively seek human assistance when needed during task execution. The ultimate goal in human-robot interaction tasks is to achieve natural, efficient, and safe interactions as robots work with humans across various domains. Mehta et al. [23] introduce a learning formalism that unifies approaches for physical human-robot interaction by incorporating demonstrations, corrections, and preferences. It represents a comprehensive approach to learning from human interactions, aiming to improve robot adaptability and performance in collaborative tasks. This framework is designed to learn without making assumptions about the tasks the human wants to teach the robot. The key insight of the paper is that physical human-robot interaction can be a rich source of information for teaching robots, and that by leveraging all available forms of interaction—kinesthetic guidance (demonstrations), adjustments to the robot’s motion (corrections), and evaluative feedback (preferences)—a more robust and flexible learning system can be developed. The authors propose a two-step algorithm that first learns a reward model from scratch by comparing the human’s input to nearby alternatives and then applies constrained optimization to map the learned reward into a robot trajectory. This process is iterative and allows for real-time updates based on the human’s feedback, which can be provided in any order and combination. The approach is validated through simulations and a user study, demonstrating that it can more accurately learn manipulation tasks from physical human interaction than existing baselines, especially when faced with new or unexpected objectives.[23] The paper’s insight emphasizes the importance of a unified learning approach that does not rely on predefined task features or reinforcement learning, thus enabling robots to learn new and unexpected tasks in real-time from physical interaction with humans. This has significant implications for the development of robots that can adapt to a wide range of tasks in shared human-robot environments, such as factories, homes, or healthcare settings, where safety and adaptability are paramount.
@@ 3.2 Key concepts and techniques
3.2.1 Behavioral cloning
Behavioral cloning (BC) is a straightforward approach that treats imitation learning as a supervised learning problem[30]. Given a dataset of state-action pairs from expert demonstrations, behavioral cloning directly learns a policy (mapping from states to actions) using regression or classification algorithms[37]. The policy is trained to minimize some loss function between the predicted and demonstrated actions on the training data. Based on the demonstration quality, BC is somewhat simple to implement since no extensive knowledge of the environmental dynamics is required. Being treated as supervised learning, a method that is very well studied, makes training BC algorithms computationally efficient.
Consider a dataset $\mathcal{D}=\{(s_{i},a_{i})\}_{i=1}^{M}$ consisting of state-action pairs collected from an expert policy $\pi^{*}$, where $s_{i}$ represents the state and $a_{i}$ the action taken by the expert in that state. The objective of behavioral cloning is to learn a policy $\hat{\pi}$ that approximates the expert policy $\pi^{*}$ as closely as possible.
The process involves, as seen in Figure 5:
@@ 3.2.1 Behavioral cloning
    For continuous action spaces, the mean squared error (MSE) loss is often used:
@@
Figure 5: Process of Behavioral cloning
@@ 3.2.2 Inverse reinforcement learning
Inverse Reinforcement Learning (IRL) is a method used to infer the reward function of an agent by observing its behavior within an environment. It assumes that we observe an agent following an unknown policy $\pi^{*}$ and we want to infer the reward function $R$ that this policy is optimizing. The problem is challenging because there are potentially many reward functions that could explain the observed behavior, making IRL an ill-posed problem. IRL is typically modeled as a Markov Decision Process (MDP) where the goal is to determine what objectives or values the agent is optimizing for, given its observed actions.
@@
- $S$ is a set of states. - $A$ is a set of actions. - $T$ is the transition probability matrix, where $T(s^{\prime}|s,a)$ gives the probability of transitioning to state $s^{\prime}$ from state $s$ after taking action $a$. - $\gamma$ is the discount factor, which determines the present value of future rewards. - $R$ is the reward function, which assigns a scalar reward to each state (or state-action pair).
@@
The general approach to IRL involves the following steps, see Figure 6:
@@
    Collecting Data: Observe the behavior of the expert agent and collect state-action trajectories.
@@
    Learning the Reward Function: Infer a reward function $R$ that would make the observed behavior appear optimal.
@@
  • Maximum Entropy IRL: This method assumes that the expert behaves in a way that maximizes entropy, meaning that among all policies that could explain the expert’s behavior, the one that is least committed to unnecessary constraints is chosen.
  • Bayesian IRL: This approach treats the reward function as a random variable and uses Bayesian methods to infer a posterior distribution over the reward function given the observed behavior.
@@ 3.2.3 Adversarial imitation learning
IRL algorithms seen previously have a high computational complexity [24][9] since they require the execution of RL in inner loops [11]. Adversarial imitation learning has been proposed as a solution to this computational challenge[12]. Generative Adversarial Imitation Learning (GAIL) is an adversarial imitation learning algorithm that uses the framework of generative adversarial networks (GANs) to directly learn a policy from expert demonstrations, without needing to first learn a reward function as in inverse reinforcement learning[12].
@@
GAIL leverages the expressive power of GANs to directly imitate expert demonstrations, without needing to recover a reward function explicitly as in inverse RL. The discriminator learns to distinguish expert data, providing a reward signal to optimize the policy to match the expert’s occupancy measure. This enables the imitation of complex behaviors from a relatively small number of demonstrations.
4 Strategies for uses in iMD-VR
Imitation learning in a fully simulated VR environment for interactive molecular dynamics offers several practical advantages over the current approach used in robotics. In molecular dynamics simulations, both training and inference can be conducted entirely within the virtual environment, eliminating the need to bridge the gap between simulation and physical reality. This approach is more practical for several reasons.
Firstly, there is consistency between training and deployment. Unlike robotics, where training occurs in VR but deployment happens in the real world, molecular simulations maintain a consistent virtual environment throughout. This eliminates the "reality gap" that often plagues robotic applications, where policies learned in simulation may not transfer perfectly to real-world scenarios.
Secondly, VR-based molecular dynamics simulations offer superior scalability and data generation capabilities. Researchers can generate vast amounts of training data quickly and efficiently, creating diverse scenarios and interactions without the physical constraints or safety concerns associated with real-world robotic systems. This is particularly valuable for exploring complex molecular systems and interactions that would be difficult or impossible to replicate in physical experiments. Furthermore, the simulated environment allows for precise control over all variables, enabling researchers to isolate specific factors and study their effects on molecular interactions. This level of control is often impossible or impractical in physical robotic setups. Researchers can manipulate individual atoms, adjust environmental conditions, and explore extreme scenarios that would be challenging or dangerous to replicate in the real world.
Cost-effectiveness is another significant advantage. Conducting both training and inference in a virtual environment significantly reduces hardware costs and eliminates the need for expensive robotic equipment. This makes the research more accessible to a broader range of institutions and researchers. Additionally, virtual simulations can be run on cloud-based systems, further reducing infrastructure costs and enabling collaborative research across different locations. Safety and repeatability are also key benefits. Virtual molecular dynamics simulations can explore potentially hazardous or extreme conditions without risking damage to physical equipment or compromising safety. Experiments can be repeated indefinitely with exact precision, facilitating rigorous scientific investigation and enabling researchers to explore a wider range of scenarios than would be possible in physical experiments.
@@ 4 Strategies for uses in iMD-VR
Imitation learning approaches like [53], [16], and learning from good samples could be applied to the task of threading a methane molecule through a carbon nanotube or the knot tying one which we explained previously. For instance, a pre-trained vision model in a transformer-like architecture, similar to VIOLA [53], could be used to detect task-driven relevant regions of the nanotube and methane molecule for precise action mapping. A hybrid approach combining behavior cloning and state cloning, inspired by HIL[16], could efficiently learn the manipulation task of threading the molecule through the nanotube or learn to tie a knot by observing expert demonstrations in VR. In similar implementations of imitation learning, researchers have identified several significant challenges, including covariate shift [29], causal misidentification [6], and the copycat problem [49]. Each of these challenges undermines the effectiveness of imitation learning algorithms, but the literature over the years has proposed innovative solutions to mitigate these issues, leading to more robust and generalizable models.
@@
Covariate shift [29], a prevalent issue where the training data distribution does not match the test data distribution, has been a focal point of concern. This mismatch leads to models that perform well on training data but fail to generalize to new, unseen environments. To combat this, interactive imitation learning (IL) techniques [45] such as DAgger [34] (Dataset Aggregation) have been developed. These methods iteratively refine the training dataset by incorporating data collected under the policy currently being learned, thus aligning the training and test distributions more closely. Furthermore, inverse reinforcement learning (IRL) approaches that focus on learning the underlying reward function from expert demonstrations offer another avenue to address covariate shift. By concentrating on the reward structure rather than directly mimicking actions, these methods aim to achieve better generalization. Additionally, constrained IL [3] introduces constraints into the learning problem to prevent significant deviations from the expert policy, thereby reducing the impact of covariate shift.
@@
Causal misidentification [6], where models learn incorrect causations between observations and actions, poses another challenge. This issue can lead to models that make decisions based on spurious correlations, resulting in suboptimal or incorrect behavior. Researchers have tackled this problem by applying causal inference techniques [46] to distinguish between causally relevant and irrelevant features. Moreover, structured imitation learning methods [35] that incorporate knowledge about the task or environment into the learning process help the model focus on the correct causal relationships, enhancing its decision-making capabilities.
@@
Madirolas et al. [22] demonstrates that the Wisdom of the Crowd (WOC) effect can be successfully applied to a visuomotor control task of tracing shapes on a touchscreen. The authors show that aggregating the trajectories from a large group of individuals (including children) produces a collective trajectory that is much more accurate than most individual ones. Specifically, the average error of the WOC trajectory was 2-5 times lower than individual trajectories. Importantly, this dramatic improvement required aggregating trajectories from different individuals, not just repeated trials from the same person. The WOC trajectory also outperformed over 99% of the individual trajectories in accuracy. This has important implications for citizen science projects that involve iMD-VR or data collection from large groups of non-expert volunteers. It further demonstrates that for certain types of tasks, aggregating the inputs from many citizen scientists, even if they are not individually highly skilled, can yield results superior to those produced by a single expert.
4.1 Potential applications of IL for iMD-VR
The use of imitation learning for iMD-VR has a range of potential applications. In this section, we identify two such domains: drug discovery and protein engineering, and material design.
4.1.1 Ligand/drug binding to protein

### v2

@@ 3.1.1 Manipulation tasks
IL has found a wide range of applications in robotics, demonstrating its versatility and effectiveness in enabling robots to perform complex tasks. The following subsections delve into three primary areas where IL has been significantly applied: manipulation tasks, locomotion and navigation, and human-robot interaction.
Manipulation tasks involve robots handling, moving, or altering the state of objects in their environment. In recent years, IL has been instrumental in teaching robots to perform such tasks with precision and adaptability[41]. For instance, VIOLA [42], a novel IL approach that was implemented and deployed into a real-life robot, outperforms state-of-the-art methods by 45.8% in success rate. This is achieved through the use of a pre-trained vision model which is put into a transformer-like architecture. The authors created a policy to detect task-driven relevant regions for action mapping. Another novel hybrid imitation learning (HIL) framework combines behaviour cloning (BC) and state cloning (SC) methods to efficiently learn manipulation tasks like pick-and-place and stacking[20]. This approach has been shown to significantly improve training efficiency and policy flexibility, demonstrating a performance improvement and faster training time compared to pure BC methods. Hua et al. [13] emphasize the efficiency of learning from good samples and the potential for combining reinforcement learning mechanisms to improve the speed and accuracy of IL. They specifically addressed the application of IL in robot manipulation by observing expert demonstrations, which can be generalized to other unseen scenarios.
@@ 3.1.2 Locomotion and navigation
In robotics, locomotion and navigation are two fundamental aspects that enable robots to move and operate within their environments effectively. These concepts are crucial for the development of autonomous systems that can perform a wide range of tasks, from simple delivery services to complex exploration missions. Locomotion refers to the various methods that robots use to move from one place to another. This movement can be achieved through different mechanisms, depending on the robot’s design and the environment it is intended to operate in. Navigation involves the process by which a robot determines its position in the environment and plans a path to reach a specific destination.
Seo et al. [43] introduce PRELUDE, a hierarchical learning framework designed to enhance the navigation and locomotion capabilities of quadrupedal robots in dynamic and cluttered environments. The framework divides the problem into two levels: high-level navigation decision-making and low-level gait generation. The high-level controller is trained using IL from human demonstrations collected with a steerable cart, enabling the robot to acquire complex navigation behaviours. The low-level gait controller is trained through reinforcement learning, allowing the discovery of versatile gait patterns through trial and error. The effectiveness of PRELUDE is demonstrated through simulations and hardware experiments, showing significant improvements over state-of-the-art reinforcement learning methods in terms of success rate and travel distance in various environmental conditions[43]. This work exemplifies the application of IL in robotics, particularly in the development of autonomous systems capable of agile and adaptive movement in real-world scenarios.
@@ 3.1.3 Human robot interaction
Human interaction tasks in robotics involve robots engaging in various forms of social interaction and cooperation with humans to achieve shared goals. These tasks encompass direct physical interaction, such as assisting with lifting objects or providing physical therapy, as well as collaborative interaction, where robots and humans work together to complete tasks like assembling products on a manufacturing line. Remote interaction, where humans control or collaborate with robots from a distance, also falls under the umbrella of human-robot interaction tasks. The ultimate goal in human-robot interaction tasks is to achieve natural, efficient, and safe interactions as robots work with humans across various domains. Mehta et al. [44] introduce a learning formalism that unifies approaches for physical human-robot interaction by incorporating demonstrations, corrections, and preferences. It represents a comprehensive approach to learning from human interactions, aiming to improve robot adaptability and performance in collaborative tasks. This framework is designed to learn without making assumptions about the tasks the human wants to teach the robot. The key insight to take away from this study by Mehta et al. is that physical human-robot interaction can be a rich source of information for teaching robots, and that by leveraging all available forms of interaction—kinesthetic guidance (demonstrations), adjustments to the robot’s motion (corrections), and evaluative feedback (preferences)—a more robust and flexible learning system can be developed. The authors propose a two-step algorithm that first learns a reward model from scratch by comparing the human’s input to nearby alternatives and then applies constrained optimization to map the learned reward into a robot trajectory. This process is iterative and allows for real-time updates based on the human’s feedback, which can be provided in any order and combination. The approach was validated through simulations and a user study, demonstrating that it can more accurately learn manipulation tasks from physical human interaction than existing baselines, especially when faced with new or unexpected objectives.[44] They emphasize the importance of a unified learning approach that does not rely on predefined task features or reinforcement learning, thus enabling robots to learn new and unexpected tasks in real-time from physical interaction with humans. This has significant implications for the development of robots that can adapt to a wide range of tasks in shared human-robot environments, such as factories, homes, or healthcare settings, where safety and adaptability are paramount.
@@ 3.2 Key concepts and techniques
3.2.1 Behavioural cloning
Behavioural cloning (BC) is a straightforward approach that treats IL as a supervised learning problem[45]. Given a dataset of state-action pairs from expert demonstrations, BC directly learns a policy (mapping from states to actions) using regression or classification algorithms[46]. The policy is trained to minimize some loss function between the predicted and demonstrated actions on the training data. Based on the demonstration quality, BC is somewhat simple to implement since no extensive knowledge of the environmental dynamics is required. Being treated as supervised learning, a method that is very well studied, makes training BC algorithms computationally efficient.
Consider a dataset $\mathcal{D}=\{(s_{i},a_{i})\}_{i=1}^{M}$ consisting of state-action pairs collected from an expert policy $\pi^{*}$, where $s_{i}$ represents the state and $a_{i}$ the action taken by the expert in that state. The objective of BC is to learn a policy $\hat{\pi}$ that approximates the expert policy $\pi^{*}$ as closely as possible.
Figure 4 illustrates the process by which this is achieved, which involves:
@@ 3.2.1 Behavioural cloning
    For continuous action spaces such as iMD-VR, the mean squared error (MSE) loss is often used:
@@
Figure 4: Process of BC
@@ 3.2.2 Inverse reinforcement learning
Inverse Reinforcement Learning (IRL) is a method used to infer the reward function of an agent by observing its behaviour within an environment [47, 48]. It assumes that we observe an agent following an unknown policy $\pi^{*}$ and we want to infer the reward function $R$ that this policy is optimizing. The problem is challenging because there are potentially many reward functions that could explain the observed behaviour, making IRL an ill-posed problem. IRL is often modelled as a Markov Decision Process (MDP) [48] where the goal is to determine what objectives or values the agent is optimizing for, given its observed actions.
@@
  • $S$ is a set of states.
  • AA is a set of actions.
  • $T$ is the transition probability matrix, where $T(s^{\prime}|s,a)$ gives the probability of transitioning to state $s^{\prime}$ from state $s$ after taking action $a$.
  • $\gamma$ is the discount factor, which determines the present value of future rewards.
  • $R$ is the reward function, which assigns a scalar reward to each state (or state-action pair).
@@
Figure  5 demonstrates the general approach to IRL which involves the following steps:
@@
    Collecting Data: Observe the behaviour of the expert agent and collect state-action trajectories.
@@
    Learning the Reward Function: Infer a reward function $R$ that would make the observed behaviour appear optimal.
@@
  • Maximum Entropy IRL: First introduced by Ziebart et al.[49] this method assumes that the expert behaves in a way that maximizes entropy, meaning that among all policies that could explain the expert’s behaviour, the one that is least committed to unnecessary constraints is chosen.
  • Bayesian IRL: This approach first introduced to IRL by Ramachandran et al.[50] treats the reward function as a random variable and uses Bayesian methods to infer a posterior distribution over the reward function given the observed behaviour.
@@ 3.2.3 Adversarial imitation learning
IRL algorithms seen previously have a high computational complexity [51][52] since they require the execution of RL in inner loops [53] (shown in Figure 5 as the IRL Algorithm box connection to Optimal Policy and meaning that for each candidate reward function, a complete RL problem to find the optimal policy is solved). Adversarial IL has been proposed as a solution to this computational challenge[53]. Generative Adversarial Imitation Learning (GAIL) is an adversarial IL algorithm that uses the framework of generative adversarial networks (GANs) to directly learn a policy from expert demonstrations, without needing to first learn a reward function as in IRL[53].
@@
GAIL leverages the expressive power of GANs to directly imitate expert demonstrations, without needing to recover a reward function explicitly as in inverse RL. The discriminator learns to distinguish expert data, providing a reward signal to optimize the policy to match the expert’s occupancy measure. This enables the imitation of complex behaviours from a relatively small number of demonstrations.
4 Strategies for applying IL to iMD-VR
iMD-VR represents a paradigm shift in how researchers interact with and manipulate molecular systems, offering an intuitive and immersive approach to exploring conformational spaces. However, the manual nature of iMD-VR interactions, while valuable for gaining physical intuition about molecular systems, poses limitations in terms of scalability and reproducibility when generating large datasets of conformational pathways. The integration of IL with iMD-VR presents a promising solution to these challenges, as it enables the capture and reproduction of expert manipulation strategies in an automated fashion. By learning from demonstrations performed by experienced researchers in VR environments, IL algorithms can extract the underlying patterns and decision-making processes that guide successful conformational transitions, effectively creating a bridge between human intuition and computational efficiency. Furthermore, In iMD-VR, both training and inference can be conducted entirely within the virtual environment, eliminating the need to bridge the gap between simulation and physical reality, which is the case in robotics. This approach is more straightforward for several reasons.
The combination of IL and iMD-VR is particularly valuable for generating comprehensive datasets of conformational pathways, as it addresses several key limitations of traditional MD approaches. First, it allows for the systematic reproduction of rare events and complex transitions that might be difficult to sample using conventional simulation methods. Second, the learned models can generalize across similar molecular systems, potentially enabling the rapid exploration of conformational spaces for novel compounds based on previously learned manipulation strategies. Furthermore, the resulting datasets can serve as valuable training resources for developing more sophisticated machine learning models, creating a positive feedback loop that continuously improves our ability to predict and understand molecular behaviour. This synergistic approach not only accelerates the generation of conformational pathway datasets but also preserves the physical intuition and expertise embedded in human-guided explorations, making it an invaluable tool for advancing our understanding of molecular systems.
@@ 4 Strategies for applying IL to iMD-VR
For instance, a pre-trained vision model in a transformer-like architecture, similar to VIOLA [42], could be used to detect task-driven relevant regions of the nanotube and methane molecule for precise action mapping. A hybrid approach combining behaviour cloning and state cloning, inspired by HIL [20], could efficiently learn the manipulation task of threading the molecule through the nanotube or learn to tie a knot by observing expert demonstrations in VR. In similar implementations of IL, researchers have identified several significant challenges, including covariate shift [55], causal misidentification [56], and the copycat problem [57]. Each of these challenges undermines the effectiveness of IL algorithms, but the literature over the years has proposed innovative solutions to mitigate these issues, leading to more robust and generalizable models.
@@
Covariate shift [55], a prevalent issue where the training data distribution does not match the test data distribution, has been a focal point of concern. This mismatch leads to models that perform well on training data but fail to generalize to new, unseen environments. To combat this, IL techniques [58] such as DAgger [59] (Dataset Aggregation) have been developed. These methods iteratively refine the training dataset by incorporating data collected under the policy currently being learned, thus aligning the training and test distributions more closely. Furthermore, IRL approaches that focus on learning the underlying reward function from expert demonstrations offer another avenue to address covariate shift. By concentrating on the reward structure rather than directly mimicking actions, these methods aim to achieve better generalization. Additionally, constrained IL [60] introduces constraints into the learning problem to prevent significant deviations from the expert policy, thereby reducing the impact of covariate shift.
@@
Causal misidentification [56], where models learn incorrect causations between observations and actions, poses another challenge. This issue can lead to models that make decisions based on spurious correlations, resulting in suboptimal or incorrect behaviour. Researchers have tackled this problem by applying causal inference techniques [61] to distinguish between causally relevant and irrelevant features. Moreover, structured IL methods [62] that incorporate knowledge about the task or environment into the learning process help the model focus on the correct causal relationships, enhancing its decision-making capabilities.
@@
4.1 Proposed approach
As an initial proof of concept we investigated the nanotube threading task as a BC challenge through a supervised learning framework, specifically employing a Convolutional Neural Network (CNN) architecture. CNNs have proven effective in dealing with spatial data where they show good performance in extracting hierarchical patterns. In our example, CNN will process spatial coordinates of all particles in the system as input, represented as 3D positional data for each simulation frame. The network’s primary objective is to predict two critical parameters: the optimal interaction position vector and the corresponding force scale factor to be applied to the central carbon atom (C61) of the methane molecule. This formulation transforms the complex threading process into a time dependant regression problem, where the model learns the mapping between the instantaneous molecular configuration and the required force parameters that would facilitate efficient passage of methane through the carbon nanotube. By leveraging the CNN’s ability to capture spatial hierarchies and local correlations in the input data, we aim to automate the threading process while maintaining physical consistency in the MD simulation. The predicted force parameters will be directly applied to guide the methane molecule’s trajectory, potentially optimizing the threading efficiency compared to traditional methods.
4.1.1 Network Design
As an initial architecture we propose a model which employs a deep CNN structure comprising five convolutional layers with increasing feature dimensionality (32 to 512 channels). Each convolutional layer utilizes a kernel size of 3 with stride 1 and appropriate padding to maintain spatial dimensions. Max pooling operations with kernel size 2 and stride 2 are applied between convolutional layers to reduce spatial dimensionality while retaining salient features progressively. The choice of this architecture is set a an initial approach mirroring previous successful approaches. In future work, we intend to explore other approaches as highlighted in literature. The architecture can be represented as:
Input(195, 1)Conv1D(1→32)(195, 32)MaxPool (97, 32)Conv1D(32→64) (97, 64)MaxPool (48, 64)Conv1D(64→128) (48, 128)MaxPool (24, 128)Conv1D(128→256) (24, 256)MaxPool (12, 256)Conv1D(256→512) (12, 512)MaxPool (6, 512)FC(3072→512) (512)FC(512→256) (256)FC(256→128) (128)FC(128→32) (32)FC(32→4) (4)Output (4)
Figure 6: CNN Architecture with Tensor Shapes
The fully connected layers progressively reduce dimensionality from 3072 to the final output dimension of 4, representing the predicted interaction parameters. ReLU activation functions are employed throughout the network to introduce non-linearity.
4.1.2 Training Methodology
The training process utilized a comprehensive dataset derived from 20 expert-guided virtual reality recordings of successful methane-nanotube threading operations whith 80% used for training and 20% for validation. Each frame in the training set was processed independently, containing a total of 65 atomic positions: 60 positions from the carbon nanotube structure and 5 positions representing the methane molecule (one central carbon and four hydrogen atoms). The target variables for each frame consisted of a four-dimensional vector, comprising the three-dimensional interaction position $(x,y,z)$ and a scalar force scale parameter. The atomic positions were arranged as a one-dimensional sequence to facilitate the CNN’s ability to extract spatial features and inter-atomic relationships. Training was conducted using the Adam optimization algorithm with an initial learning rate of 0.001, employing Mean Squared Error (MSE) as the loss function to minimize prediction errors. Model performance was evaluated using the R² score metric, which quantifies the proportion of variance in the target variables that is predictable from the atomic positions. This frame-by-frame approach ensures that the model learns the temporal independence of the threading process while capturing the essential spatial correlations between atomic positions.
4.2 Initial Results
(a) CNN Training Loss
(b) CNN Validation Loss
Figure 7: CNN loss evolution plots
Figure 8: R2 score during training
Figures 7 and 8 show that despite its architectural simplicity, this CNN implementation demonstrates promising results in predicting user interactions from particle positions alone.The training performance analysis of our CNN model for methane-nanotube threading reveals a comprehensive learning progression over 5000 training steps. The model demonstrates distinct learning behaviour throughout the training process, starting with negative R² values around -0.2, indicating initial predictions below baseline performance. The learning trajectory shows remarkable improvement in the early stages, with a particularly steep ascent during the first 2000 steps. During this period, the R² values rapidly improve from negative territory to approximately 0.8, demonstrating the model’s growing capability to capture the underlying physics of methane-nanotube interactions. As training progresses beyond the 2000-step mark, the learning curve exhibits a more gradual improvement, characteristic of fine-tuning behaviour. The model ultimately achieves a final R² score of approximately 0.87, indicating that it successfully explains 87% of the variance in the interaction position and scale predictions. This high performance level suggests excellent capture of the essential features governing methane-nanotube threading dynamics. The training curve’s stability is particularly noteworthy, displaying consistent behaviour with minimal variance throughout the learning process. The narrow confidence band surrounding the main training curve indicates robust feature extraction. This stability, combined with the high final performance metrics, suggests that the model has successfully learned to predict the complex interactions between methane molecules and carbon nanotubes, making it a valuable tool for MD applications. Several avenues for architectural enhancement should be explored:
  • Integration of physics-informed constraints in the network design
  • Incorporation of temporal dependencies between consecutive frames
  • Extension to graph-based architectures to better capture molecular topology
  • Implementation of attention mechanisms for focusing on relevant atomic interactions
These enhancements could lead to more robust and generalizable models for predicting molecular interactions in virtual reality simulations, potentially improving the accuracy and efficiency of MD studies.
5 Discussion and future directions
In this paper our results show that an agent’s ”intelligence” can be derived from an IL model trained on datasets generated by human experts within iMD-VR environment. When embodied as an avatar and reintegrated into the iMD-VR space, the agent’s primary function is currently to clone the experts’ behaviour by mimicking their actions which it successfully does. Judging by the current state of literature and the advancements made to IL and iMD-VR, and by combining these approaches with MAS IL approaches the agent’s function may evolve beyond simple task replication. The focus shifts to the dynamics of human-agent interaction, creating a collaborative partner for research or training. For a teaching system, this avatar would not only demonstrate expertise by generating viable pathways for molecular interactions but would also need to engage with the human user in a natural and responsive manner. The success of this integration hinges on the agent’s ability to model the complex, nuanced behaviours of a dyadic interaction, making the virtual collaboration feel seamless and intuitive, thereby effectively augmenting the human user’s capabilities. In this section, we identify two possible domains for the application of IL for iMD-VR: ligand/drug binding to proteins and material properties investigation. We also discuss the prospect of employing citizen science approaches to create datasets for training IL models.
5.1 Ligand/drug binding to proteins